the demand function of good1 is estimated as q1=1000-150p+2y+20p2 if current prices good 1 is 5 birr the price of good 2 is 8 birr and par capital of income is 500 birr where q1is quantity of good 1 demand, p1 is prices of good 1, p2 is price of good 2,and y is per capita income of the consumer then calculate 1.prices elasticity of demand and elasticity type 2.cross prices elasticity of demand and the relation between the two products 3.income elasticity of demand and state the type of product
1
Expert's answer
2021-08-20T00:45:01-0400

1.price elasticity demand

= change in quantity demaded/change in price

in equation;

q1=1000-150(5) + 2(500) + 20(8)

q1= 1410

q2 = 100-150(8) + 2(500) + 20(5)

q2= 900

PED = 1400/900 /(5/8)

=2.5

the demand is elastic.


2.cross price elasticity demand

Exy= Change in Qx/Qx/change in Py/py

0.36/-0.6

-0.6

the goods are compliments.


2.income elasticity demand

= change in quadity demanded/ change in capital income

=1410/900/(1000/500)

= 0.78

They are normal goods.